The December 31, 2015, partial financial statements taken from the annual report for AT&T Inc. (T ) follow.

Consolidated Statements of Income
Dollars in millions except per share amounts 2015 2014
Operating revenues
Service $ 131,677 $ 118,437
Equipment 15,124 14,010
Total operating revenues 146,801 132,447
Operating expenses
Equipment 19,268 18,946
Broadcast, programming and operations 11,996 4,075
Other cost of services (exclusive of depreciation and amortization show separately below) 35,782 37,124
Selling, general and administrative 32,954 39,697
Abandonment of network assets -- 2,120
Depreciation and amortization 22,016 18,273
Total operating expenses 122,016 120,235
Operating income 24,785 12,212
Other income (expense):
Interest expense (4,120) (3,613)
Equity in net income of affiliates 79 175
Other income (expense) - net (52) 1,581
Total other income (expense) (4,093) (1,857)
Income before income taxes 20,692 10,355
Income tax expense 7,005 3,619
Net income $ 13,687 $ 6,736
Consolidated Balance Sheets -- Liabilities and Equity Sections
Dollars in millions except per share amounts, December 31 2015 2014
Current liabilities
Debt maturing within one year $ 7,636 $ 6,056
Accounts payable and accrued liabilities 30,372 23,592
Advanced billed and customer deposits 4,682 4,105
Accrued taxes 2,176 1,091
Dividends payable 2,950 2,438
Total current liabilities 47,816 37,282
Long-term debt 118,515 75,778
Deferred credits and other noncurrent liabilities:
Deferred income taxes 56,181 38,436
Post employment benefit obligation 34,262 37,079
Other noncurrent liabilities 22,258 17,989
Total deferred credits and other noncurrent liabilities 112,701 93,504
Stockholders' equity
Common stock ($1 par value, 14,000,000,000 authorized atDecember 31, 2015 and 2014; issued 6,495,231,088 atDecember 31, 2015 and 2014) 6,495 6,495
Additional paid-in capital 89,763 91,108
Retained earnings 33,671 31,081
Treasury stock (350,291,239 at December 31, 2015 and1,308,318,131 at December 31, 2014, at cost) (12,592) (47,029)
Accumulated other comprehensive income 5,334 8,061
Noncontrolling interest 969 554
Total stockholders' equity 123,640 90,270
Total liabilities and stockholders' equity $ 402,672

$296,834

Consolidated Statements of Stockholders' Equity -- Excerpts 2015
Amount in millions except per share amounts, December 31 Shares Amounts
Common Stock
Balance at beginning of year 6,495 $ 6,495
Issuance of stock -- --
Balance at end of year 6,495 $ 6,495
Additional Paid-In-Capital
Balance at beginning of year $ 91,108
Issuance of treasury stock (1,597)
Share-based payments 252
Change related to acquisition of interests held by noncontrolling owners --
Balance at end of year $ 89,763
Retained Earnings
Balance at beginning of year $31,081
Net income attributable to AT&T ($2.37, $1.24 and $3.42 per diluted share)) 13,345
Dividends to stockholders ($1.89, $1.85 and $1.81 per share) (10,755)
Balance at end of year $ 33,671
Treasury stock
Balance at beginning of year (1,308) $(47,029)
Repurchase of common stock (8) (278)
Issuance of treasury stock 966 34,715
Balance at end of year (350) $(12,592)

(a) How much interest expense did AT&T incur during 2015?
(b) What is the book value of AT&T's interest-bearing debt at the end of 2015?
At the beginning of 2015?
Average debt for 2015? (Round to the nearest whole number)

(c) Estimate AT&T's 2015 pretax cost of debt capital.

Round your answer to one decimal place.

(d) Estimate AT&T's 2015 effective (that is, average) tax rate from information in its income statement.

Round your answer to one decimal place.

(e) Using your rounded answer from (c) above, estimate AT&T's 2015 after-tax cost of debt capital. The company's statutory tax rate is: 37%.

Round your answer to one decimal place.
Answer #1

(a). Interest expense for 2015 is $ 4120 million.

(b). Interest bearing debt means , the amout which company borrowed as short term and/or long term debt it can be from finnacial institiutions, from Hire purchase, from related party and from anyone one on which we paid int against that amount.

In this scenario in 2015 begining we have a debt of ($ 6,056 + $ 75,778) = $ 81,834 milion and

At the end of 2015 we have a debt of ($ 7,636 + $ 118,515) = $ 126,151 million

Average debt = (opening debt + closing debt) / 2 = (2015 + 2016) / 2 which is ($ 81,834 + $ 126,151) / 2 = $ 200,349 / 2 = $ 100,174.50 million

(c) . Pre-tax cost of debt is also sometime referenced as the effective interest rate. it's not widely used, since the effective interest paid is tax deductible.

Pre Tax cost of debt capital = Total Interest payments / Total outstanding debt

Total interest Paid in 2015 = $ 4,120 million.

Total Outstanding debt = ( $ 7,636 + $ 1,18,515 ) = $1,26,151 million.

Cost of Debt = 4,120 / 1,26,151 = 3.3% (Approx)

(d). Average tax rate = Total income tax paid / Income before income tax.

Total income tax paid during 2015 = $ 7,005 million.

Total income before income tax for 2015 = $ 20,692 million.

Average rate of tax for 2015 = 7,005 / 20,692 = 33.9 %

(e). Cost of debt = 3.3 % (From point C )

Rate of tax given = 37 %

Post tax cost of debt capital = (1 - Tax rate ) * Pre tax cost of debt.

= (1 - 0.37 ) * 0.033 = 0.02079 = 2.1%